OpenStudy (anonymous):

In a school, 2/5 of pupils are boys and the rest are girls. If there are 165 girls, how many boys are there?

OpenStudy (campbell_st):

start with 3/5 = 165 then find 1/5 1/5 = 165/3 1/5 = 55 then double to find the number of boys

OpenStudy (anonymous):

110

OpenStudy (campbell_st):

yes.. thats it

OpenStudy (anonymous):

Herrro

OpenStudy (anonymous):

your such a hero

hero (hero):

Am I?

OpenStudy (anonymous):

Yeah hero

OpenStudy (anonymous):

yes you are

OpenStudy (anonymous):

your brown?